Beta Company purchased machinery on June 1, 2018 for $89,000. It provided the following data:
Estimated Residual Value - $5 000 Estimated Service Life – 8 years
Estimated Total Work Hours – 42 000 hours Estimated Total Production – 525 000 units Machine Usage (2018) – 5 500 hrs
(2019) – 6 000 hrs Machine Production (2018) – 48 000 units (2019) – 55 000 units

Beta Company uses a fractional-year policy to the nearest full month and its financial year ends December 31.
a) What is the depreciable base of the machine?

b)Calculate the depreciation expense for the year ending December 31, 2018 and 2019
using the following methods:
Straight-line.
i.Units of output.
ii.Sum of the years’ digits.
iii.Double declining balance.
iv.Round to nearest whole number
Answer #1

(a)

depreciable base = initial cost - salvage value

= $89000 - $5000

= $84000

(b)

(1)

Under straight line method of depreciation,

Depreciation per annum = (initial cost - salvage value)/useful life

= ($89000 - $5000)/8

= $10500

As annual depreciation expense is uniform each year,

for the year ending December 31, 2018

Depreciation expenses = $10500

for the year ending December 31, 2019

Depreciation expenses = $10500

(2)

Under units of production method,

Depreciation per annum = (initial cost - residual value) x (units produced during the year/estimated total production units)

Therefore,

for the year ending December 31, 2018

Depreciation = ($89000 - $5000) x (48000/525000) = $7680

for the year ending December 31, 2019

Depreciation = ($89000 - $5000) x (55000/525000) = $8800

(3)

Under sum-of-the-years'-digit method,

Depreciation per annum = (initial cost - salvage value) x (remaining useful life of asset/sum of the years' digit)

Sum of the years' digit = 1+2+3+4+5+6+7+8 = 36

Therefore,

for the year ending December 31, 2018

Depreciation expenses = ($89000 - $5000) x (8/36) = $18,666.67

for the year ending December 31, 2019

Depreciation expenses = ($89000 - $5000) x (7/36) = $16,333.33

(4)

Under double declining balance method,

Depreciation per annum = 2 x straight line depreciation percentage x book value at the beginning of period

And,

Straight line depreciation percentage = 100%/useful life = 100%/8 = 12.5%

Therefore,

for the year ending December 31, 2018

Book value in the beginning = $89000

Depreciation expenses = 2 x 12.5% x $89000 = $22250

Book value at the end = $66750

for the year ending December 31, 2019

Book value in the beginning = $66750

Depreciation expenses = 2 x 12.5% x $66750 = $16,687.5